OpenStudy (anonymous):

Arjun spent Rs.12 to purchase some pens and pencils.The cost of a pen and pencil are in the ratio 3:2. the total cost of a pen and a pencil is Rs. 3. if the total number of pens and pencils he purchased is7, find the number of pens he purchased.

OpenStudy (anonymous):

@AravindG

OpenStudy (shubhamsrg):

What was your attempt ? Did you try ?

OpenStudy (aravindg):

ya this one is too easy ..can u show the equations you wrote?

OpenStudy (anonymous):

Rs 3 = 3 : 2 (pen to pencil) Rs 3 = 3x+2x Rs 3 = 5x Rs 3/5 = x = Rs 0.60 or paise 60 Therefore one pen and pencil together cost Rs0.60 * 5 = Rs 3.00 Therefore cost of 7 pen and 7 pencil = Rs 15.00

OpenStudy (anonymous):

plse see if i have done correct or not

OpenStudy (shubhamsrg):

How do you figure 15 ? Seems wrong to me .

OpenStudy (shubhamsrg):

the questions tells you he purchased some pens and pencils, and asks you what no. of pens he purchased if total cost was 7